### v3

# InstructAny2Pix: Flexible Visual Editing via Multimodal Instruction Following

Abstract
The ability to provide fine-grained control for generating and editing visual imagery has profound implications for computer vision and its applications. Previous works have explored extending controllability in two directions: instruction tuning with text-based prompts and multi-modal conditioning. However, these works make one or more unnatural assumptions on the number and/or type of modality inputs used to express controllability. We propose InstructAny2Pix, a flexible multi-modal instruction-following system that enables users to edit an input image using instructions involving audio, images, and text. InstructAny2Pix  consists of three building blocks that facilitate this capability: a multi-modal encoder that encodes different modalities such as images and audio into a unified latent space, a diffusion model that learns to decode representations in this latent space into images, and a multi-modal LLM that can understand instructions involving multiple images and audio pieces and generate a conditional embedding of the desired output, which can be used by the diffusion decoder. Additionally, to facilitate training efficiency and improve generation quality, we include an additional refinement module. These designs are critical to the performance of our system. We demonstrate that our system can perform a series of novel instruction-guided editing tasks.
Keywords:
Image Editing Multi-Modal Text-to-Image
1 Introduction
Figure 1: The InstructAny2Pix pipeline consists of three building blocks: a multi-modal encoder that encodes different modalities into a unified latent space, an LLM that can encode instructions (encoding [gen]), and a diffusion model that learns to decode latent representations into edited images. For improved training and generation, we include an additional prior module to refine the LLM outputs.
With growing fidelity of generative models of images [4, 24], the ability to control their outputs is critical for many real-world applications, ranging from creative generation to synthetic augmentations. However, popular state-of-the-art tools such as ControlNet[37] and T2I[22] can only perform specific edits they are trained on. Many of these models also require inputs such as Canny edge, Depth, or Surface Normal, making them inaccessible to general users without expertise in computer vision. To mitigate this, instruction-based image editing methods such as InstructPix2Pix[5] allow users to describe their instructions in natural languages, such as “add a dog." However, such methods are still limited to simple instruction on which they are trained and cannot generalize to complex instructions involving multiple editing operations or multiple objects. Additionally, their inability to take additional audiovisual inputs, makes them struggle with tasks such as style transfer.
We propose InstructAny2Pix, the first instruction-following image editing system that can follow complicated, multi-modal, multi-object instructions. Concrete examples of such instructions can be “add the [sound] to [image]," where the sound can be that of a dog barking or a piece of music. It can also be “add [object A] and remove [object B] from [image]," where objects can be represented through either images, text, or audio. Additionally, Our system also accommodates free-form instructions like “change [image A] to the style of [image B]" or “fit [image] to [music]." This work significantly broadens the scope of image editing instructions. Through both quantitative and qualitative evaluations, we show that our proposed method achieves high performance on diverse editing tasks.
Our framework brings together the multi-modal perception capability of a multi-modal encoder, high quality generation capability of a diffusion model and instruction understanding capability of a LLM. The performance of our system depends on three key components:
First, we train our model on large amounts of multi-modal edit instructions. Our training data consists of diverse instructions, including incorporating audio into an image, adding an image to an image, adding multiple objects, or replacing an object with another object where objects are represented by either images, text, or audio. We generate this dataset by prompting state-of-the-art Large Language Models with human-crafted examples for each category of instructions. This diverse dataset is crucial for extending instruction-following capabilities of image-editing systems to long, complex, and multi-modal instructions.
Second, we integrate recent techniques in instruction tuning of language models and multi-modal representation learning to efficiently understand instructions involving inputs of multiple modalities. In particular, we encode the multi-modal inputs using the encoder and insert them into the instruction sequence that is passed to a LLM. The LLM then learns to understand the instruction and generate an output embedding. This embedding is then used to condition a diffusion model that generates the image.
Last, we propose a refinement prior module to facilitate efficient training and enhance generation quality. The need for such a refinement module arises from two key observations. First, large language models are slow to train, and it may take considerable time for them to learn to generate high-quality representations that are useful for generation tasks. Second, the quality of data typically used for multi-modal alignment is worse than that used to train high-quality generative models. For example, the average aesthetic score for VGG-Sound [6], a dataset commonly used for audiovisual alignment, is only 4.5 because it mostly consists of low-quality YouTube clips. In contrast, LAION-600M [28], which is typically used to train diffusion models, has a minimum score of 5. These observations make it necessary to introduce a method to mitigate unwanted biases and accelerate training. We implement our refinement prior module as a transformer that learns to refine the output embedding of the LLM.
Figure 2: Illustration of InstructAny2Pix’s ability to flexibly edit an image based on a variety of different multi-modal instructions.
2 Related Works
2.1 Image Editing
There are a large number of image-editing methods based on text-to-image diffusion models [2, 26, 24, 12]. They can be generally categorized into three families.
The first family is based on Image2Image translation. Works like ControlNet [37] and T2I [22] added convolution adaptors to the U-Net of Diffusion Models. These models enable image generation conditioned on texts and additional inputs such as canny edge, and depth. They achieve image editing by first translating the image to the desired conditional domain and generating a new image from this condition using desired text prompts. However, such methods lose considerable information from the source image and struggle to preserve details due to the translation mechanism. These methods also have poor accessibility due to the amount of expertise required to generate these conditions.
The second family is based on text prompts. These methods only require prompts describing the desired output, making them more accessible. The naive approach is to perform DDIM[29] inversion that converts an image back into the latent space of a diffusion model and generates a new image from such a latent representation. Prompt2Prompt (P2P) [11] improves this baseline by injecting the cross-attention maps during the diffusion process. Plug-and-Play [32] additionally injects convolution features to provide more refined control. Parallel to these lines of work, Null-text Inversion [21] proposed a better inversion method for input images through a learnable null text prompts. While these methods are easier to use than the previous family, they still require long detailed descriptions of each desired output. This limits their usability in applications like batched editing.
The last family is instruction-guided methods. Unlike previous methods, these models only requires a vague text instruction such as “add fireworks". InstructPix2Pix [5] achieves this by curating a large machine-generated image editing dataset using P2P, then directly fine-tuning on instruction-following generation. MagicBrush [36] curated a higher quality human-annotated dataset by requesting humans to perform editing operations using tools such as PhotoShop. MGIE [7] utilizes a multi-modal large language model to process editing instructions and input images. But it still operate on a single source image with text-only instructions. All these works are limited to simple instructions and require iteratively applying the model recurrently for complex instructions. Additionally, they cannot take audiovisual input as auxiliary information.
Unlike previous works in this area, our work extends the scope of valid instructions to multi-modal, multi-object inputs and can perform complicated instructions in a single turn.
2.2 Multi-Modal Conditioned Generation
Parallel to these image editing methods, there are some previous attempts to achieve image generation with multi-modal conditioning using Multi-modal language models. BLIP-Diffusion [15] incorporates BLIP [16] as a multi-modal encoder that generates subject embedding for the diffusion model. Using this approach, it was able to generate images that following text prompts and reference images. Kosmos-G [23] also makes use of multi-modal language models. By directly aligning the representation space of multi-modal language models with that of a diffusion model, Kosmos-G allows image generation following multiple reference images. However, since these works focus on generation rather than instruction-based image editing, they requires compositional text inputs. These typically follow patterns like “[x] in [y] wearing [z]". They do not support removal and replacement of object, and general free-form instructions They also cannot faithfully respect the spatial structure of reference images.
Audio guided image generation is a relatively uncharted area. AAI [34] achieves sound-guided generation by aligning audio representation to reference images. This method is very limited in that it requires retrieving 3-5 reference images and performing gradient descent optimization steps for each audio input at inference time.
Unlike previous works in this area, our work is the first to support interleaved audiovisual inputs and free-form image editing instructions.
Figure 3: Data generation Pipeline. 1. We prompt LLM with sampled captions and examples to generate a diverse set of instructions. 2. We generate a candidate dataset using a variety of caption or segmentation based methods, and filter them into a final dataset using CLIP score. 3. We show some example image pairs in the filtered dataset.
3 Methods
Figure 4: Training pipeline of InstructAny2Pix. 1.Language Alignment 2. Generation Learning 3. Refinement Learning 4. Realigning Diffusion Model 5.Instruction Finetuning
Our pipeline is illustrated in Figure 1. To generate a sample of an edited image given a multi-modal instruction, we first leverage a pretrained multi-modal encoder to encode non-textual inputs into a set of embeddings. We then pass these embeddings alongside the text instruction to a large language model and generate a sequence of control tokens. Each control token belongs to one of the following type: [base], [gen], [sub]. The embedding of the [base] token is used to retrieve the source image of editing operation. The embedding of [sub] tokens is used to retrieve reference images of relevant subjects in the desired output. Typically there is only one [base] token, but there can be more than one [sub] tokens to account for multiple reference images. The embedding of [gen] tokens is used to condition the diffusion model. It will be further refined by the refinement module. Finally, the diffusion model generates an image output from retrieved source image, the retrieved embeddings of reference images and refined embeddings of [gen] tokens.
The training pipeline of InstructAny2Pix  consists of five stages: Multi-Modal Language Alignment, Multi-Modal Generation Learning, Refinement Learning, Realigning Diffusion Models and Instruction Guided Finetuning.
3.1 Multi-Modal Language Alignment
We use a MLP projector to map the output of multi-modal encoder to the embedding space of LLM. The goal of multi-modal alignment stage is to learn a good mapping between these two spaces. We make use of a large number of image-text pairs and audio-text pairs. We freeze all parts of the model other than the projector. We prompt the model with the text instruction “describe the [x]", where [x] is the embedding of an image or audio piece. We use the ground truth caption as the target sequence. Standard cross-entropy loss is applied to the output of the LLM.
3.2 Multi-Modal Generation Learning
We use a MLP projector to map the output of LLM back to the representation space of the multi-modal encoder. For clarity, we refer this projector as the output projector, and the projector described in Sec. 3.1 as the input projector. We freeze all parts of the model other than the input and output projector. We make use of image-text pairs and image-audio pairs in this stage. We prompt the model with instruction “generate [x]", where [x] is either a ground truth caption or an embedding of an audio piece. We use the ground truth image embedding encoded by the multi-modal encoder as the target, and apply L2 loss on the output projector. Since the final linear head of LLM is frozen, we cannot obtain [gen] token in the output. We assume the [gen] sequence comes immediately after user input, and directly retrieve the last-layer embedding of LLM at corresponding location as inputs to the output projector.
3.3 Refinement Learning
We observe that there are many low quality images used in the alignment step, most of which come from audio-image pairs. Particularly, many of these pairs come from internet videos with low resolution, bad exposure, undesired motion blur, and other artifacts. We empirically verified this observation by evaluating the LAION aesthetic scores of the dataset. To mitigate the effect of low quality data, we incorporate a refinement module, which is a transformer that learns to improve the image quality in the embedding space. Since pairs consisting of one image and its improved version is hard to obtain, we make use existing image-text and image-audio pairs. As the embedding space of the multi-modal encoder is aligned, we can directly model $\text{Pr}(x|e,s)$, where $x$ is the image, $e$ is the embedding and $s$ is aesthetic score. We can use the text and audio embedding during the training, and substitute them with image embedding at inference.
In this stage, only the refinement module is used. Given image-text and image-audio pairs, the refinement module takes the aesthetic score of the image and the embedding of text or audio as the input, and learns to predict the embedding of the ground-truth image. We apply the L2 regression loss on the predictions.
3.4 Realigning Diffusion Models
The goal of this stage is to realign a text-to-image diffusion model to accept conditional embeddings in the representation space of multi-modal encoder. In particular, it should learn to generate image from both refined [gen] embeddings which encode some global semantics and reference image embeddings retrieved by [sub] token. We employ an MLP projector to map the embeddings to the representation space of the diffusion model. In this stage, we train U-Net of the diffusion model and the MLP projector, keeping the VAE of the diffusion model unchanged. We only use high quality image filtered by aesthetic score in this stage and train the model in a self-supervised fashion. We use the ground truth image embedding as [gen] input, and image embedding of a cropped region as [sub] input.
There are several challenges: first, there may be a distribution shift between the ground truth image embedding and those generated by the refinement module; second, in the training process the [gen] and [sub] input comes from the ground truth image, hence it is important to prevent the model from overfitting to the global embedding and ignore the region embedding. To address the first challenge, we add random gaussian noise to [gen] embedding to improve the robustness. To address the second challenge, we randomly drop the global and/or region embedding independently at a probability of 0.5. Additionally, we randomly perform unconditioned inpainting in the cropped region so that [gen] embedding comes from an image with the region corresponding to [sub] repainted. Because inpainting is expensive, we pre-select a subset and generate these images offline.
In this stage, we apply the standard diffusion loss as [26]. The regions used to crop the image are either generated by an object detector or randomly sampled at 1:1 ratio. The regions used for inpainting are all generated by detector.
3.5 Instruction Guided Finetuning
In this stage, we fine tune the LLM and projectors end-to-end on multi-modal edit instructions and pairs of source image and edit results. We defer the details of data generation process to Sec. 3.7. The input projector is used to map the encoded embeddings of reference images and audio pieces to the representation space of LLM, while the output projector is used to map the LLM output back to the representation space of multi-modal encoder. We formulate the target output sequence as “[base] [sub] .. [sub] [gen] .. [gen]" where each [sub] corresponding to a reference image. We only add a “[sub]" token if the referenced object should appear in the desired output. For example, in the following instructions “add [a]", “remove [b]", “replace [c] with [d]", only [a] and [d] will have corresponding [sub] tokens in the target sequence. We apply cross entropy classification loss on the output of the final linear head of LLM, as the standard language training. Additionally, we fetch the corresponding embeddings of target sequence from the last hidden layer of LLM and project them into the representation space of multi-modal encoder through output projector. L2 regression loss is applied between the [base] embedding and ground truth embedding of the source image, between [sub] embeddings and ground truth embeddings of reference images, and between the [gen] embedding and ground truth embeddings of the target image. The diffusion model is not used in this process. We do not need to actually generate images as the loss is directly applied in the embedding space.
3.6 Implementation Details
We use ImageBind [10] as the encoder and SDXL[24] as the decoder. We use Vicuna-7B[30] as the base LLM. We use two-layer MLP with GELU activation for the input and output projectors.
3.7 Datasets
3.8 Paired Dataset
For image-text pairs, we use 2.4M text-image pairs sampled from LAION-Aesthetic-V2 datasets. Because the vanilla captions are of low quality, we augment the captions by recaptioning the images using BLIP2 [16]. For audio-text pairs, we make use of AudioCaps[13] and MusicCaps [1] dataset. For audio-image pairs, we make use of 2M audio-visual pairs sampled from VGG-sound [6], AudioSet [8], and SoundNet [3]. These dataset allow InstructAny2Pix  to learn a broad range of textual-visual and audio-visual correspondences.
3.9 Instruction Tuning Dataset
For instruction tuning, we curated a diverse dataset of 500k instructions and corresponding image pairs, called MM-Inst. The dataset generation pipeline consists of two steps: instruction generation and image generation.
In the instruction generation phase, we prompted a Large Language Model (LLAMA2 [31]) to generate creative instructions using 36 manually written examples. The example instructions include adding, dropping, removing, or replacing objects as well as other free-form instructions. To generate a diverse set of image editing instructions,we prompt the language model to create editing instructions based on sampled captions from LAION dataset. To further increase the diversity of audio-related instructions, we further prompt the language model to generate instructions involving ground truth music captions from MusicCaps and AudioCaps. We observe that without this step, the language model tends to only generate simple audio captions such as “sound of water" or “sound of rain" and fails to incorporate complex descriptions of music. In total, we generate 500k instructions. We provide further details in the appendix.
In the image generation phase, we generate corresponding image pairs of editing instructions. Unlike previous works such as InrstructPix2Pix [5] which only uses off-the-shelf caption guided image editing methods like Prompt2Prompt [11], we make use of combination of following techniques.
  • Edit image using captions and Prompt2Prompt [11]
  • Edit image using captions and Plug-and-Play [32]
  • Use DDIM[29] inversion on the source image and generate a new image using target prompt with inversed latent.
  • For object removal, we use an open-vocabulary object detector (GroundingDINO [19]) to locate the object and perform inpainting with remove the object.
  • For object addition, we first generate an image using target prompt as the target image. Then we use the detector to localize the added object and remove it through inpainting. The resulting image is used as the source image.
  • For object replacement, we first follow the removal procedure. Then we perform inpainting on the area of removed objects using the replacement object as prompts.
When the detector fails to localize the object or output low confidence scores, we fall back to the caption-based methods. Following InstructPix2Pix [5] we filter the results using CLIP scores. We additionally filter the results using the LAION aesthetic predictor and remove low quality images. We show this pipeline in Fig. 3
3.10 Evaluation Dataset
Image editing following multi-modal instruction is a novel task with no existing benchmarks. To fairly evaluate InstructAny2Pix’s performance in real-world settings, we curated 1500 manually written multi-modal instructions. We employ Amazon Mechanic Turk for this task. To generate a diverse set of prompts, we ask the MTurk workers to generate creative edit instructions using different captions sampled from LAION. We also require the MTurk workers to generate different types of edits for the same caption. We do not generate ground truth target images. For reference images and audio pieces used in the instructions, we use SDXL [24] to generate images and AudioLDM2 to generate [17] audio. Results are filtered by CLIP[25] score for images and CLAP[33] score for audios. Notable, unlike previous works which performs evaluations on samples from the same distribution of training data, our evaluation benchmark caters for diverse real-world usecases. We call this dataset MM-Inst-Test.
For image guided generation, Dreambooth[27] is a common benchmark used. However we find it inadequate to provide a holistic evaluation of the capability of multi-modal generations. Particularly, there are two limitations. Firstly, its prompts consists only two classes of live objects (cats and dogs), which in total has 9 subjects out of 30. Hence, it has very limited diversity. Secondly, its task only involve changing the background of a single subject. This setup cannot evaluate the capability of models capable of understanding multiple images. To address this gap, we propose Dreambooth++, a Dreambench dataset of more diverse prompts. It consists of 30 subject images which are evenly distributed across humans, animals, small objects, and large structures. We also includes 30 diverse background images with corresponding prompts. In total, there are 900 generation tasks. We evaluate models on this dataset using two protocols: single-image and multi-image. The single image setup is similar to Dreambooth, which requires generating a given subject under different context prompts. The multi-image task requires generating a new image by combining a subject image and a background image. We defer further details to the appendix.
Additionally, for completeness and fair comparison with existing models, we evaluate our model on 1000 samples from InstructPix2Pix dataset. Note that we do not train our model on this dataset, and the evaluation is performed in a zero-shot manner.
4 Experiments
Table 1: Multi-Modal Image Editing on MM-Inst-Test Dataset and Text-based Image Editing on InstructP2P Dataset. (I+T+A) and (T) refers to using multi-modal instruction and text-only instruction respectively.
MM-Inst InstructP2P
CLIPdir CLIPim CLIPout Win. CLIPdir CLIPim CLIPout Zero-shot
Ours(I+T+A) .099 .816 .260 -
Ours(T) .095 .856 .270 - .147 .808 .312
InstructP2P .091 .824 .243 .712 .145 .742 .241
MagicBrush .084 .807 .199 .707 .165 .760 .250
InstructDiff. .066 .940 .193 .746 .126 .857 .301
Table 2: Image Conditioned Generation on DreamBench++ Dataset. $C_{dir},C_{im},C_{out}$ is abbreviated form of CLIPdir, CLIPim and CLIPout. For multi-image setup, numbers are reported in DINOref/DINOsub format.
Single-Image Multi-Image
Cdir Cim Cout DINO Cdir Cim Cout DINO Recall
Ours(T+I) .147 .810 .260 .688 .154 .789 .309 .625/.471 .841
BLIP-Diffusion .089 .779 .231 .660 .091 .701 .292 .526/.422 .693
Kosmos-G .126 .843 .251 .683 .166 .740 .286 .485/.476 .812
Figure 5: Music-Gudied Image Variation: Music uniquely conveys emotions that are hard to describe using other modalities such as language. We recommend the reader find the corresponding music on Apple Music or other platform for optimal experience.
Figure 6: Ablation Studies of hyperparameters.
(a) Ablation Study on the effect of hyper-parameters and the refinement module. Increasing the cfg and decreasing the alpha will increase CLIPout (CLIP-T) and decrease CLIPim (CLIP-I). Adding refinement module improves both metircs.
(b) Ablation Study on Image Consistency. Notably, our method generate artifact-free results in all setup, while other methods have visiable artifact at all cfgs.
Figure 7: Multi-object Multi-modal Instruction Editing. Unlike previous methods that rely on an iterative chain of editing to perform complex instructions, our model can directly understand complex instructions involving multiple subjects of multiple modalities and perform them in a single turn. For comparison, we also show results of applying each edit sequentially.
We qualitatively and quantitatively evaluate our results on a wide range of challenging instructions with varying combinations of multi-modal inputs.
4.1 Instruction guided Image Editing
[…]
We also perform human evaluations. Human evaluators are asked to pick the preferred edit output in one-to-one comparison between $InstructAny2Pix$ and baseline methods. For a fair comparison, we use the text-only version of our method. We report win rate in Tab. 1.
We show decisive advantages in human preference and strong performance in quantitative evaluation metrics comparing with baseline methods InstructPix2Pix[5], Magicbrush [36] and InstructDiffusion [9]. Notably, we achieve competitive performance on InstructPix2Pix dataset without ever training on such dataset. This result showcases the superiority of our data generation pipeline. It incorporates a diverse range of instructions and enables our model to generalize to unseen instruction patterns.
4.2 Image Conditioned Generation
We evaluate our method on DreamBench++ dataset described in Sec. 3.10. We conduct both single-image and multi-image evaluation and compare results against Blip-Diffusion [15] and Kosmos-G [23]. In addition to metrics reported in the previous section, we additionally report DINO score which measures the image similarity. For multi-image benchmark, evaluating the DINO similarity of the entire image does not make sense as the subject is added to the scene and not necessarily occupies the entire image. To address this, we use a segmentation model [19] to segment the subject in generated images. We report the DINO similarity between the cropped subject against the referecne subject image as DINOsub, and the DINO similarity between the whole result and the background input image as DINOref. We provide quantitative metrics in Tab. 2 and qualitative comparison in the appendix. . InstructAny2Pix  shows a clear advantage in generation quality and image consistency.
4.3 Multi-Object Multi-Modal Instruction Editing
Unlike previous works such as [5], our methods is capable of performing complex editing operations involving many inputs of different modalities. Compared with previous methods, our method drastically extends the boundary of instruction-guided image editing.
4.4 Ablations
4.4.1 Refinement
To study the effectiveness of our refinement method, we compare the editing results with refinement and without refinement. We find that it leads to improvement in edit quality measured by CLIPim (+.007) and CLIPout (+.03). We also observe significant improvements in visual quality of outputs. We provide additional studies in the appendix.
4.4.2 Control Strength
An important goal of image editing is to ensure the edited images can reflect the intended changes while respecting the source image. There is usually a trade-off between these two goals. The relevant hyper parameter of InstructAny2Pix  is cfg and $\alpha$. CFG is the strength of classifier-free-guidance in the generation process. It determines the degree at which the control signal affects the generation output. We visualize the results in Fig. 6(b). We find that CFG=5 is a sweet spot for achieving high quality edit results that follows the instructions while respecting the original image. To make sure our model can follow the spatial structure of reference image, we don’t initialize the latent in the generation process from a canonical normal distribution. Instead, we sample a latent from a gaussian distribution centered at the DDIM inverse of the source image, with a variance proportional to $\alpha$. We define further details to appendix. Intuitively, $\alpha$ is the hyperparameter that trades image consistency with generation diversity. We visualize the effect of CFG and $\alpha$ in Fig. 6(a). Generally, increasing the cfg and decreasing the alpha will increase CLIPout and decrease CLIPim. For a typical use, we suggest setting $\alpha$ to 1.0.
5 Discussion
5.1 Quality of Generation
We observed that our method generates higher quality outputs compared with previous works that also make use of SDXL [24]. We hypothesize this may be caused by our data isolation. In particular, we only train on high-quality images in a self-supervised manner, so the model does not inherit bias from editing datasets that have a lower average quality than the dataset used to pretrain SDXL. While we made this choice primarily for training efficiency, we hope future work can explore the corruption in quality of diffusion models when fine-tuned for editing.
5.2 Music Guided Generation
InstructAny2Pixuniquely enables music guided image editing, which can be quite useful in cases like automatic album generation using portraits of artists. To evaluate this potential, we perform music-guided image variation where we prompt the model with the instruction "modify [image] to convey the feeling of [music]". We find the model can successfully adapt an image to different music in subtle ways hard to be described in other modalities.
6 Conclusion
In summary, we propose InstructAny2Pix, a flexible system for editing images based on multi-modal, multi-object instructions. Compared with previous works, we significantly expand the scope of instructions and are capable of performing complex instructions without resorting to a recurrent chain-of-editing. To efficiently train our system which contains an LLM with billions of parameters, we proposed a novel, decoupled training scheme that significantly improved the training efficiency by removing end-to-end fine-tuning. We hope future works can better address the scalability of the overall system when coupling many modality specific encoders and large language models.
Appendix
7 Additional Experiment Results
7.1 Qualitative Comparison with Text-based Methods
We provide a qualitative comparison with methods using text instructions. We compare our results against InstructPix2Pix [5], MagicBrush [36], and Instruct Diffusion [9] in Fig. 8. For fairness, we incorporate instructions from both the InstructPix2Pix dataset and the MM-Inst-Test dataset. For MM-Inst-Test, we convert the multimodal instructions to text by replacing the multimodal token with captions of the referred image and audio. Our model shows better editing results on both datasets. This is impressive since our model is not trained on the InstructPix2Pix dataset while other methods are.
Figure 8: Qualitative Comparison Against Text-based Editing Methods. We show editing results of different models using diverse editing instructions. The top three rows are sampled from the InstructPix2Pix dataset, and the bottom three rows are sampled from the MM-Inst-Test dataset.
7.2 Qualitative Comparison with Image-conditioned Generation Methods
We provide a qualitative comparison with multi-modal generation methods that use images as references. We compare our results against BLIP-Diffusion [20] and Kosmos-G [23] in Fig. 9. Visual results show that InstructAny2Pix  outperforms these two baselines both in terms of generation quality and image consistency. We also conducted human evaluation using Amazon Mechanical Turk. We asked users to pick the best result in a side-by-side comparison of InstructAny2Pix  and baseline methods using generations of Dreambooth++ dataset. We achieved a win rate of 79.0% against BLIP-Diffusion and 86.2% against Kosmos-G.
Figure 9: Qualitative Comparison Against Image-based Generation Methods. We present generation results of different models under both single-image and multi-image setups. We employ multi-modal prompts from the Dreambooth++ dataset.
7.3 Additional Ablation Studies
7.3.1 Image Quality
We present qualitative results of multi-modal editing in Fig. 10. The results demonstrate that InstructAny2Pix  can effectively handle a diverse range of instructions involving multiple modal inputs.
Figure 10: Additional Results of Multi-Modal Editing. We showcase qualitative results of multi-modal editing. InstructAny2Pix  can handle a diverse set of instructions involving multiple modalities.
7.3.2 Image Consistency
In the main paper, we provided quantitative analysis of $\alpha$ and CFG in Sec. 4.4.2. We provide further qualitative results in Fig. 11. We compare generations at different $\alpha$ with an empty edit instruction. Recall that $\alpha$ controls the noise added to the latents obtained through DDIM inverse. When $\alpha=1$, no noise is added to inversed latents, leading to generations that closely follw the source image spatially.
Figure 11: Results of Varying $\alpha$ in the generation process. We visualize the image generation conditioned on the CLIP embedding of the source image. As $\alpha$ decreases, the generation become less consistent with the source image.
7.4 Anonymous Demo Website with Audio
Due to the nature of this work, we have created an anonymous demo website with playable audio at https://sites.google.com/view/annodemo3333/project-page. Since some music belongs to respective copyright owners, we have embedded publicly available YouTube videos instead of hosting the audio files ourselves. We are not the uploaders of these YouTube videos, and therefore they do not contain any identifiable information.
7.5 Image Quality Ablation
We sample 500 captions and evaluate the Aesthetic score and PickScore [14] for samples generated with and without the refinement process. These two metrics are known to reflect the visual quality of generated images. Aesthetic scores only consider the image quality, while PickScore additionally takes the prompt into account. On both metrics, adding the refinement module leads to considerable improvements. We show these results in Fig. 12.
Figure 12: Quantitative Ablation results of Refinement Module. We report average Aesthetic score and PickScore [14] on 500 randomly sampled captions. Aesthetic scores only consider the image quality, while PickScore additionally takes the prompt into account.
8 Additional Details about Datasets
8.1 Paired Training Data
We use SoundNet [3], VGG-Sound [6], and AudioSet [8] for image-audio pairs. These datasets consist of videos with audio. We extract the audio and the middle frame from the video to create audio-image pairs. SoundNet consists of 802,724 audio-image pairs, AudioSet consists of 2 million audio-image pairs, and VGG-Sound consists of 197,958 pairs. Particularly, out of 2 million videos in AudioSet, 1 million are under the music category. These videos can be music videos, concerts, documentaries, and other kinds of videos that use music as background, such as news programs. They naturally allow InstructAny2Pix  to learn a generic correspondence on how humans associate images and music. The rest of the audios are mostly sounds of objects. They help InstructAny2Pix  to associate the visual appearance of an object to its sound. We highlight this in Fig. 13.
We also make use of audio captions from MusicCaps [1] and AudioCaps [13] to create text-audio pairs. These two datasets provide text captions for subsets of AudioSet. They do not introduce new audio files. We use LAION-Aesthetic-3M [28] for text-image alignment, which consists of 2,209,745 valid image URLs at the time of data fetching (Sep 2023). All these datasets are used in the alignment process.
Figure 13: Example Thumbnails of Videos with Music. We show samples of thumbnails of videos from the music category of the AudioSet dataset. It incorporates diverse music-image correspondences.
8.2 MM-Inst
8.2.1 Implementation Details
In this section, we provide additional details of the generation process of multi-modal inputs. For audio and music, there are two types of inputs. The first type is generated by explicitly prompting the LLM with ground truth captions of music or sound. In this case, the corresponding audio can be directly used. The second type is created solely by the LLM. We use AudioLDM2 [18] to generate 5 samples for each of these captions and use CLAP [33] to find the sample that aligns best with the caption. For image inputs, we generate the image using SDXL from captions. For image pairs generated specifically through the segmentation+inpaint pipeline, we directly crop the images using the bounding boxes of objects and use the results as multi-modal inputs.
8.2.2 Comparison with Prior Works
Compared with the InstructPix2Pix dataset, MM-Inst has the following advantages. First, it uses BLIP-generated captions which are grounded with real images instead of the raw caption from LAION, which can be very noisy. Second, it uses a variety of techniques to generate paired data instead of solely relying on Prompt2Prompt. In particular, we observe that the segmentation+inpainting pipeline generates high-quality results for object removal when the segmentation model can correctly localize the target object. Third, we filter the data using both CLIP and Aesthetic score and consider both prompt alignment and generation quality. In contrast, InstructPix2Pix only uses CLIP score as the filtering mechanism. Lastly, MM-Inst incorporates multi-modal inputs, which makes it uniquely suitable for our multi-modal editing tasks. We showcase these differences in Tab. 3. We provide examples from both datasets in Tab. 4.
MM-Inst MM-Inst-Test InstructPix2Pix
Caption Source LAION-Aesthetics LAION-Aesthetics LAION-Aesthetics
Input Caption Generation BLIP2 BLIP2 Noisy WebData
Instruction Generation Llama 2 Human GPT-3
Paired Data Generation DDIM - Prompt2Prompt
Prompt2Prompt -
Plug-and-Play -
Segmentation+Inpaint -
Filtering Multiple Metrics Human CLIP
Modality Image,Text,Audio Image,Text,Audio Text
Size 500,000 1,500 313,010
Table 3: Comparison of MM-Inst dataset and Instruct Pix2Pix dataset.
MM-Inst
Instruction Output
Please incorporate [an image of cannon fire] into [an image of a pirate ship sailing on the high sea] An image of a pirate ship firing at a British Navy warship, fire burning on the ship
Remove [sound of car accelerating] from [an image of people driving in the countryside road] An image of a quiet countryside road
Replace [sound of dog barking] with [sound of a cute cat] for [an image of a dog at the beach] An image of a cat at the beach
Change [an image of a woman wearing sunglasses in Paris] to the style of [an image of a Renaissance painting of a noble lady] A Renaissance painting of a woman wearing sunglasses in Paris
Make [an image of a cute girl in a school uniform] fit the atmosphere of [a piece of music of stellar constellations] An image of a cute girl in a school uniform under the night sky
InstructPix2Pix
[misurina XIII… by roblfc1892] have it be a stamp Stamp… misurina XIII… by roblfc1892
[Manarola during sunrise - Cinque Terre] it is foggy Manarola during foggy sunrise - Cinque Terre
Table 4: Examples of Instructions from MM-Inst dataset and Instruct Pix2Pix dataset. Captions are marked by [.]. MM-Inst offers a better set of captions and instructions.
8.3 MM-Inst-Test
One of the challenges in generating the test dataset is that there are instances of bad formatting and low quality. We prompt each MTurk worker to generate five different edit prompts for each caption. After a batch is collected, we manually identify the problematic instructions and redistribute them to a new set of workers. In order to reach 1,500 valid instructions, we collect a total of 1795 instructions. We show the distribution of different types of instructions in table Fig. 14.
Figure 14: Distribution of Instructions in MM-Inst-Test dataset. We show the number of instructions. "Fit" refers to combining music and image, fitting an image to the style of another image, or other organic ways of combining different modalities together. "Others" include all instructions that cannot be classified as other categories, such as "transform [image] into a night scene with the sound of [sound]".
8.4 Dreambooth++
Dreambooth is commonly used to evaluate generation with image prompts; however, it only contains multiple images of the same subject, not multiple images of different objects. Moreover, the classes have limited diversity and only have two live subjects (cats and dogs). To provide a diverse and fair comparison, we propose Dreambooth++, which contains 30 subjects and 30 prompts. In total, there are 900 combinations. We use SDXL to generate all subject images. For each prompt, we also generate a background image.
We conduct evaluation on two benchmarks: The single-image setup is similar to Dreambooth, which requires generating a given subject under different context prompts. The multi-image task requires generating a new image by combining a subject image and a background image. For the multi-image task, we use a segmentation model to localize the edited objects in the new image. We report DINO similarity of the cropped subject with the reference subject image as DINOsub and the DINO similarity of the generated image with the reference background image as DINOref. We also report the recall rate of the segmentation model. These results are listed in the main paper Tab. 2.
Additionally, we provide a full list of subjects in Tab. 5 and compare it with Dreambooth. DreamBooth++ offers a more diverse range of subjects ranging from animals, humans, large structures, and small items.
Table 5: Subjects of DreamBooth++ and DreamBooth Dataset. DreamBooth++ offers a more diverse range of subjects ranging from animals, humans, large structures, and small items.
DreamBooth++ Dataset Subjects DreamBooth Subjects
a cute cat backpack
a cute dog backpack_dog
a colorful butterfly bear_plushie
a colorful bird flying low over a body of water berry_bowl
spotted horse can
an image of a squirrel candle
an african elephant walking through a grassy field cat
an image of cute anime girl cat2
an anime princess with long blonde hair and two swords clock
the woman in a black dress holding a fan colorful_sneaker
a man standing at a podium dog
a cyberpunk style dog2
an image of a scientist dog3
an image of an astronaut dog5
an artistic painting of a woman with blonde hair dog6
a wooden bridge dog7
car dog8
traffic lights duck_toy
train fancy_boot
tree grey_sloth_plushie
bicycle monster_toy
an image of a robot pink_sunglasses
tablet poop_emoji
telescope rc_car
the skull red_cartoon
vase robot_toy
wand shiny_sneaker
an image of chair teapot
an image of empty glass vase
a cowboy hat wolf_plushie
9 Implementation Details
9.1 Multi-Modal Encoder
We use ImageBind [10] as our multi-modal encoder. Particularly, ImageBind uses CLIP-ViT-L as its text and image encoder. It includes an additional music encoder that is aligned to the representation space of CLIP-ViT-L. We use the pooled token as our multi-modal embedding.
9.2 Diffusion Model
We use the SDXL [24] as our diffusion model. It was originally conditioned on CLIP-ViT-G and CLIP-ViT-L text features. We incorporate an MLP projection layer following [35] that maps the ImageBind embedding to the dimension of cross-attention layers. During the training process, the loss is
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t}\left[\lVert\epsilon_{t}-p_{\theta}(z_{t},C_{t},C_{g},C_{l})\rVert_{2}^{2}\right]$ (1)
where $p_{\theta}$ is the U-Net, $\epsilon_{t}$ is the noise at timestamp $t$, $z_{t}$ is the noised version of the image in latent space at time $t$, $C_{t}$ is the CLIP embedding of captions. $C_{g}$ is the embedding of the whole image, $C_{l}$ is the embedding of a cropped region. The cropped region is sampled from SAM masks or a uniform distribution of bounding boxes at a 1:1 ratio. At each training step, $C_{t}$, $C_{g}$, $C_{l}$ are randomly dropped independently with a probability of 0.2.
9.3 Refinement
We adopted a decoder-only transformer with 24 layers and a hidden size of 1024. We also incorporate an MLP projector that maps ImageBind features to the hidden dimension of the transformer. Another MLP projector is used to map the output of the transformer back to the dimension of ImageBind features.
9.4 Multi-Modal Large Language Model
We use Vincuna-7B [30] as our base model. We made no additional changes to the LLM architecture other than adding input and output projectors described in Sec. 3.
9.5 Parameter Count
We report the total number of parameters in each module in Tab. 6. In total, our model has around 10B parameters.
Table 6: Number of Parameters in InstructAny2Pix. We report the total number of parameters in each module.
Params
LLM 7B
SDXL 2.5B
Refinement 71.1M
Projectors 62.9M
10 Limitations
10.1 Dataset
Our model makes use of pretrained a diffusion model [24] and LLM [30]. Hence, it may inherit biases from the training process of these models.
10.2 Generation Style
Our model tends to bias towards artistic/painting outputs instead of photorealistic ones. This is caused by multiple factors: First, the LAION-Aesthetic-3M [28] dataset used to recondition the diffusion model contains a lot of art and paintings. Additionally, the LAION Aesthetic score used to condition the prior model is biased towards high saturation and artistic outputs. Lastly, we use SDXL to generate images for the MM-Inst dataset based on captions. Without explicit style keywords in prompts, we find that SDXL generations are biased towards artistic outputs as well. We will try addressing this limitation by exploring alternative ways of curating a high-quality dataset and explicitly adding diverse style prompts in the generation process.